SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

   
Prev Next


Date: February 28, 2008
Shift Taker: Graham Jackson
-----------
General Features:
(theta,phi)

Warm crystals in hitmap	(52,63), (56,63), (51,64)
Warm crystals in 0.01 GeV hitmap (17,43), (30,45), (32,46), (36,64)
Warm crystals in 0.1 GeV hitmap (30,45), (32,46), (36,64)
Warm crystals in 1 GeV hitmap      

Cool crystal  in hitmap          
Cool crystal  in 0.1 GeV hitmap (29,84) 
Cool crystal  in 1 GeV hitmap    		

[1] =  TC Damage:OOS/OOS ~10^-3 in 0-5 all fibers. (And usually FLinkA/B/C Timeout is high too).
[2] =  PEP were having problems according to the log book.
         The energy scale seems odd on the 1GeV hit map!?...Too warm?
         Large peaks in theta, phi projections suggest high bgs. Low digi multiplicity above the 1GeV cut.
         Low level noise throughout the digi timing plot and also (sometimes) higher level noise the the left of the peak, AND a small 'blob' to the rhs. 

0079996 [1]. [2].
0079997 [1]. [2].
0079998 [1]. [2].
0079999 [2].
0080000 [2].
0080001 ok.
0080002 [1].
0080003 [1].
0080004 ok.
0080005 [1]. Evidence of noisy channels around the peak.
0080006 Evidence of noisy channels around the peak.
0080007 [2].
0080009 ok. Short run.
0080010 [2]. Short run.
0080012 [2].  V odd 1GeV hit map. PEP had just lost the LER.
0080013 [1].
0080014 TC Damage, OOS at 10^-3: on 1-3 all fibers.
0080015 ok.
0080016 [1].
0080017 Extremely short run.  A blob and lots of noise to the rhs of the digi timing peak.
                TC Damage, OOS at 10^-1: on 1-3 all fibers.
0080024 Cosmics, no damage.
0080025 Cosmics, no damage.
0080027 Cosmics, no damage.
0080028 Cosmics,  TC Damage, OOS at 10^-3: on 1-3 all fibers.
0080029 ok.


Crystal Spectra:
Badness 1 & 2 where ADB 3-4-1 has just been replaced.

Lightpulser:
4 channels dead, 1 channels out of time. Updated on: 28-February 2008 16:44 PST


 

  Links

  • Shift Report Index
  • Back to the Shifter instructions